consider the vector space of R3

2.1. Show that <x, y>=x1y1+2x2y2+x3y3, x="\begin{bmatrix}
 x1 \\
 x2 \\
x3
\end{bmatrix}" ,y="\begin{bmatrix}
 y1 \\
 y2 \\
 y3
\end{bmatrix}""\in" R3

 2.2. Are the vectors "\begin{bmatrix}
 1 \\
 1 \\
 1
\end{bmatrix}", "\begin{bmatrix}
 1 \\
 1 \\
 - 1
\end{bmatrix}", "\begin{bmatrix}
 1 \\
 -1 \\
 -1
\end{bmatrix}" Linearly independent?

2.3. Apply the Grem-Schmidt process to the following subset of R3 "\begin{bmatrix}
 1 \\
 1 \\
 1
\end{bmatrix}", "\begin{bmatrix}
 1 \\
 1 \\
 - 1
\end{bmatrix}", "\begin{bmatrix}
 1 \\
 -1 \\
 -1
\end{bmatrix}" to find an orthogonal basis with the inner product defined in 2.1. for the span of this subset


1
Expert's answer
2021-10-19T17:20:02-0400

2.

"\begin{vmatrix}
 1 & 1& 1\\
 1 & 1&-1\\
1&-1&-1
\end{vmatrix}=-2-2=-4"

Since determinant is not equal zero, the vectors are linearly independent.


3.

orthogonal basis:

"v_1=u_1=(1,1,1)"

"v_2=u_2-\frac{\langle u_1,u_2\rangle}{|u_1|^2}u_1=(1,1,-1)-\frac{2}{3}(1,1,1)=(1/3,1/3,-5/3)"

"v_3=u_3-\frac{\langle u_1,u_3\rangle}{|u_1|^2}u_1-\frac{\langle u_3,v_2\rangle}{|v_2|^2}v_2=(1,-1,-1)+\frac{2}{3}(1,1,1)-\frac{4}{9}(1/3,1/3,-5/3)="

"=(5\/3,-1\/3,11\/27)"


1.

Let "(e_1,e_2,e_3)" be a basis.

Then:

"x=x_1e_1+x_2e_2+x_3e_3"

"y=y_1e_1+y_2e_2+y_3e_3"


"\langle x,y \rangle =\sum_i\sum_j x_iy_i\langle e_i,e_j\rangle"

if "\langle e_1,e_1\rangle=1,\langle e_1,e_2\rangle=\langle e_1,e_3\rangle=\langle e_2,e_3\rangle=0,\langle e_2,e_2\rangle=2,\langle e_3,e_3\rangle=1"

then "\langle x,y \rangle=" x1y1+2x2y2+x3y3